OpenStudy (666potatoes):

Simplify: 8√32a^2b^5 c & √12a^2 / √3a Guide me through these please ._.

OpenStudy (anonymous):

\[8\sqrt{32a^2b^5c}\] \[8\sqrt{(16)(2)a^2b^4bc}\] \[8(4)ab^2\sqrt{2bc}\] \[32ab^2\sqrt{2bc}\]

OpenStudy (666potatoes):

Thank you thank you. c:

OpenStudy (anonymous):

\[\frac{ \sqrt{12a^2} }{ \sqrt{3a} }\] \[\frac{ \sqrt{12a^2} }{ \sqrt{3a} }\times \frac{ \sqrt{3a} }{ \sqrt{3a} }\] \[\frac{ \sqrt{36a^aa} }{ 3a }\] \[\frac{ 6a \sqrt{a} }{ 3a }\] \[2\sqrt{a}\]